OpenStudy (anonymous):

The measurements of a chemical storage container are shown in meters. Find the cost of painting the exterior of nine of these large cylindrical containers with sealant. The sealant costs $32 per gallon. Each gallon covers 18 square meters. Do not paint the bottom faces.

OpenStudy (matt101):

Picture?

OpenStudy (anonymous):

oh yhea oops. the diameter is 7 in. and the height is 10 in

OpenStudy (matt101):

You wrote 18 m^2 in the original question but you gave the diameter and height in inches - is this how it is in the original question? Not a big deal, just involves an extra step for the unit conversion.

OpenStudy (anonymous):

i ment meters -.- sorry

OpenStudy (matt101):

Haha good to know... Alright well first calculate the surface area of one cylinder. The circular faces each have a surface area of pi*3.5^2 = 38.5 m^2 The side of the cylinder has a surface area of 10*pi*7 = 219.9 m^2 Since we're ignoring the bottom face, we're effectively dealing with 38.5 + 219.9 = 258.4 m^2 per cylinder. For 9 cylinders, that's 258.4*9 = 2325.6 m^2 total that have to be covered. To find the number of gallons required, divide by 18: 2325.6/18 = 129.2 gallons. Since each gallon costs $32, the total cost is 129.2*32 = $4134.4

OpenStudy (anonymous):

okay! so i guess the books answer is wrong -.- it says its 4160 ? thanks!